# Utilizing and extending superconducting circuit toolbox to simulate analog quantum gravity

### v2

@@ III Setup
Crucially, the nearest-neighbour model provides an intriguing loop hole: here, the overtilt at $k\approx 0$ has no partner point with inverted tilt, very simply because the dispersion relation satisfies the periodicity constraint in $k$ by taking, in a sense, a “detour” in the complex plane – by courtesy of the exceptional point. Consequently, the spectrum crosses the zero energy line only once. We are thus able to conclude that there is after all a case, where black and white hole horizons can be created independently in a lattice. This comes at the cost of a spontaneous instability within the overtilted region, which, importantly, is present even in the absence of an event horizon. This finding will be of great importance when interpreting the origin of radiation in the presence of event horizons in Sec. V below.
IV Realizing negative inductance through quench
@@
As noted above, an overtilt in the dispersion relation is realized by means of either negative capacitances or negative inductances. We here present ideas for the engineering of both features, and then argue, why we expect the latter to be more feasible, thus demonstrating claim (ii) in Sec. II.2,
@@
Negative (and other nonlinear) capacitances are an interesting topic within quantum circuits, which date back to various pioneering ideas by Little [106] and Landauer [107], and have recently been a focal point of research, either in the form of ferroelectric materials [108, 109, 110, 111], or capacitively-coupled polarizers inducing pairing in quantum dots [112, 113]. Specifically within the circuit QED context, it has recently been proposed that negative capacitances exist in the sense of an emf-induced renormalization [69], or that the charge in the quantum phase slip energy contribution can be fractionalized and thus rendered incommensurate [114, 115]. There is, however, the problem that negative capacitances correspond to an electrostatic instability, and are thus only meaningful either as partial capacitances (where another capacitance in parallel guarantees total positivity on a given charge island) or in a transient regime. Since we are interested in the latter (after all, Hawking radiation is manifestly a transient feature), this means that the capacitance should have to be tunable on very fast time scales, which seems challenging, especially on the small scales of the here proposed circuit QED architecture.
@@
This is why we move on to ideas on how to bring about a transient behaviour within the inductive (Josephson junction) element. Remember that we focus on a regime of $E_{J}>E_{C}$ where the cosine of the Josephson energy can be approximated as a parabola with the inductive energy $E_{L}\sim E_{J}$ of an effectively regular linear inductor (rendering the resulting plasmon field effectively non-interacting, as shown above). Nonetheless, we can exploit the nonlinear inductive behaviour. Suppose there is a mechanism, allowing for a fast, transient switch from $-\cos(\phi_{j}-\phi_{j-1})$ (e.g., for the nearest neighbor model) to $+\cos(\phi_{j}-\phi_{j-1})$. If that switch occurs sufficiently fast, the circuit degrees of freedom have no time to adjust, such that the many-body wave function immediately after the switch remains localized on what is now the maximum (and not the minimum) of the cosine. Expanding for small phase differences localized around said maximum, we can realize a sign switch of the inductive energy, $E_{L}\rightarrow-E_{L}$, leading to the sought-after overtilt in the dispersion relation (see the orange lines in Fig. 2). Specifically, in the next section, we discuss the possibility of applying an inverted inductance only in part of the system to create event horizons separating regions with $u<|v|$ from unquenched regions with $u>|v|$.
@@
FIG. 4: Illustration of non-local effects of the current loop flux drive. In order to create the target phase shift profile of exactly $\pi$ within a connected region in the junction array (from $j=j_{0}$ to $j=j_{1}$), the corresponding applied current profile as a function of site index $j$ needs to be slightly nonlocal. We have plotted three current profiles, that were obtained for a system with 50 sites with $j_{0}=15$ and $j_{1}=35$, with different ratios of $\Delta l$ to $\Delta x$ as shown in the graph. The parameter $R$ does not have a significant effect on the nature of the current profile, only on the magnitude of the ration $I_{j}/I_{0}$, hence we set $R/\Delta x=1$.
@@
[…]
@@
For the matrix given in Eq. (23), we can explicitly compute a required current profile that needs to be applied in order to swap the phases by ±$\pm\pi$ within a connected part of the 1D chain. Blue dots in Fig. 4 show that if $\Delta l\ll\Delta x$, $\alpha$ is almost perfectly diagonal, such that there is a one-to-one correspondence between the desired target phase profile and the required current profile that has to be applied to the individual loops. Note however that the onset of non-local effects may play a role even for (moderately) small $\Delta l/\Delta x$ (orange and brown dots in Fig. 4), where we need to apply a slightly nonlocal current profile to have a sharp, local switch from positive to negative inductances within the 1D chain. To conclude this section, we point out similarities and differences to previous proposals creating analog event horizons and Hawking radiation in solid state systems. Specifically, the analog Hawking radiation previously predicted in Josephson junction arrays [32, 33] relies on the creation of solitons, where within the finite extension of the soliton (over many lattice sites), the inductance can likewise be regarded as effectively negative within the soliton profile. The procedure we produce on the other hand involves creating ±$\pm\pi$ shifts on a connected series of junctions, which in essence corresponds to a tightly packed generation of “half”-solitons with the width of a single lattice site. Our system thus seems arguably much less stable. But first, as indicated already, the nonreciprocity provided by the gyrators allows for a stabilization of the system even with inverted inductors, due to an extra (positive) inductive contribution. Second, the extra tilting mechanism provided by the gyrators allows for the explicit realization of arbitrary (on-demand) spacetime geometries, and in addition, the explicit creation of wormholes with (in principle) distinguishable black and white hole horizons. As for the evolution of the unstable system beyond transient times, we provide some thoughts in Sec. VI. Another important difference is that we here have in principle perfect control on the position of the horizon, contrary to the solitons studied in Ref. [32, 33], which are autonomously moving with respect to the lab frame. Moreover, we can create an event horizon with the precision of a single lattice site. This is in contrast to all other analog Hawking radiation proposals (cold atoms, circuits, tilted Weyl semimetals), where the analog metric always changes over a finite width (a healing length). This is an interesting caveat specifically regarding the existence of a well-defined Hawking temperature, or more generally, how to make sense of horizons with diverging curvature, as already indicated in Sec. II.
V analog horizons via flux quench
@@
Once the optimization problem of the current profile has been solved, the sign of the Josephson energy can be flipped for a region of the circuits described by the Hamiltonians (III) and (III), resulting in two horizons in the circuits (Fig. 1c). In the presence of horizons, the energy spectrum is in general complex for both nearest and next-to-nearest neighbour implementations.
@@
The presence of tilted dispersion relations and horizons can be probed by means of injecting wave packets within the chains, and probe their time of flight. For example, within an overtilted region, wave packets can only move in one direction. At a white hole horizon, an incoming wave packet comes to a stand still. Such features represent a first experimentally available signature of the nontrivial spacetime geometry. Note that for the time evolution of the wave packets, the distinction between nearest and next-to-nearest neighbour coupling is not that important, under the condition that the experimenter is able to prepare wave packets close to $k\approx 0$ or $k\approx\pi$. These two types of excitations clearly distinguish between black and white hole horizon as they probe only a local part of $k$ space. In addition, aspects of the wormhole stability (or instability) are not immediately visible in the wave packet amplitude (until the moment when the evaporation starts changing the spacetime geometry itself, see also Sec. VI below).
@@
We therefore present in addition a more sophisticated measurement which is able to probe radiation due to instabilities in a more direct fashion. Namely, this section examines how to characterize the system via two point charge and phase correlation functions. As foreshadowed above, while for a stable system, quantum fluctuations remain bounded over time, they start to diverge in the presence of complex eigenvalues. In the quadratic approximation, see Eq. (III), it is convenient to express the charge and phase operators for the circuits (III) and (III), in terms of bosonic operators
@@
that obey the following commutation relations $[a_{j},a^{\dagger}_{j^{\prime}}]=\delta_{jj^{\prime}}$, $[a_{j},a_{j^{\prime}}]=0$. Now, calculating the two point correlation functions for charge and phase operators turns into calculating the correlation functions of the form $\langle a_{j}(t)a_{j^{\prime}}^{\dagger}(t^{\prime})\rangle$, $\langle a_{j}^{\dagger}(t)a_{j^{\prime}}^{\dagger}(t^{\prime})\rangle$, $\langle a_{j}(t)a_{j^{\prime}}(t^{\prime})\rangle$, and $\langle a_{j}^{\dagger}(t)a_{j^{\prime}}(t^{\prime})\rangle$.
We are going to calculate these correlations using two methods: 1) direct diagonalization and 2) an extension of Klich’s determinant formula [120]. The reason for the multi-pronged approach will be explained in detail below. To summarize, 1) is in principle very straightforward to implement and program, and is more versatile, as it would allow computing all types of observables, not only correlations. But diagonalization is not applicable in the presence of exceptional points, which is where method 2) comes into play, since it only requires exponentiation of the matrix with exceptional points.
As previously stated, once the system has been quenched, its state is in a highly excited state, were the new ground state is removed from the theory (a well-defined procedure for times sufficiently close to the quench). In order to anchor the theory to a well-defined ground state (necessary to compute observables in a well-defined way), we choose to perform computations with respect to the ground state of the unquenched system. Since the diagonalization of the quenched Hamiltonian is being performed with respect to a state that is in general not its eigenstate, we will need a generalised version of bosonic Bogoliubov transformations. These transformations are discussed in Appendix B.
Implementing the Bogoliubov transformations on the quenched Hamiltonian, mentioned in the previous paragraph, requires diagonalizing a non-Hermitian matrix. This can pose a problem if the matrix has defective eigenvalues (exceptional points). We therefore suggest another method of calculating the correlations after the quench. Consider the following way to write a two point correlation
$\left\langle a^{\dagger}_{i}(t)a_{j}(t)\right\rangle=\partial_{\chi}\left\langle e^{\chi a^{\dagger}_{i}(t)a_{j}(t)}\right\rangle_{\chi=0}$
$=\lim_{\beta\to\infty}\frac{\text{Tr}\left(e^{i\mathcal{H}_{>}t}e^{\chi a^{\dagger}_{i}a_{j}}e^{-i\mathcal{H}_{>}t}e^{-\beta\mathcal{H}_{<}}\right)}{\text{Tr}\left(e^{-\beta\mathcal{H}_{<}}\right)},$ (26)
@@
The system we will study here has fifty sites ($J=50$) with the wormhole located between $j_{0}=15$ and $j_{1}=35$. For convenience, we impose periodic boundaries in position space (i.e., the 1D chain is closed into a loop). The parameters of the system are $|E^{\prime}_{L}|=|E_{L}|=3.42E_{C}$, $G=2.4$ and $M=3*10^{-4}E_{C}$. As already pointed out, in each of the following figures of correlation functions the result could be obtained by both diagonalization and the generalized Klich determinant. The time steps in the following plots are $\delta t=0.031E_{C}^{-1}$ for the circuit with nearest neighbor coupling and $\delta t=0.31E_{C}^{-1}$ for the next-to-nearest neighbor coupling.
FIG. 5: System with Josephson junctions connecting the nearest neighbors, (left) time evolution of quantum fluctuations of phase difference, (right) time evolution of quantum fluctuations of conjugate charge. Red dotted lines denote the position of the apparent horizons, i.e., boundaries between the wormhole and normal regions. Here, the fluctuations visibly distinguish between a pure black (pure white) hole horizon, where quantum fluctuations accumulate (or not). More over, the interior between the two horizons is here unstable, such that quantum fluctuations diverge immediately within the entire wormhole region.
FIG. 6: System with Josephson junctions connecting the next-nearest neighbors, (left) time evolution of quantum fluctuations of phase difference, (right) time evolution of quantum fluctuations of conjugate charge. Red dotted lines denote the position of the horizons. Here, (and contrary to Fig. 5) both horizons have black as well as white hole character, such that the two horizons are not easily distinguised when considering the spatial dependence of the fluctuations. On the other hand, the wormhole interior is here (at least initially) stable, and quantum fluctuations grow only from the horizons outwards.
@@
First, let us examine the plots of correlation functions for the circuit with only nearest neighbor coupling, Fig. 5. We see that both phase and charge quantum fluctuations diverge after the quench. The system has two boundaries (red dotted lines) between the wormhole and the normal regions, that act as the horizons (located at $j_{0}$ and $j_{1}$). If we initialize low energy excitations (for this circuit this means excitations near $k=0$) in the region with overtilted dispersion, they move from the left horizon to the right horizon, hence we label them as the black and the white hole horizon respectively. The black hole horizon radiates away with time, as can be inferred by the accumulation of quantum fluctuations near it with time (which indicates presence of radiation), while the white hole horizon does not radiate. This explicitly confirms the previously discussed fact that for the nearest neighbour system, black and white hole horizons are distinguishable. The wormhole region also starts radiating due to the presence of the exceptional point (as likewise already discussed in Sec. III). We observe that the decay in the wormhole interior is much faster than the one induced by the presence of the horizons, which leads to the correlation functions growing much more rapidly inside the wormhole, before the black hole horizon show any appreciable decay. Overall, the above thus illustrates the possibility to create lattice simulations of wormholes with distinguishable black and white hole horizons, at the cost of a strong instability of the wormhole interior.
@@
Now, let us focus on the circuit where Josephson junctions connect next nearest neighbors. Low-energy excitations in this circuit can be either near $k=0$ or $k=\pm\pi$. For excitations (inside the wormhole region) near $k=0$, similar to the previous circuit, the left boundary ($j_{0}$) acts as the black hole horizon and the right one ($j_{1}$) as the white hole. In contrast, for excitations near $k=\pm\pi$ the left boundary acts as the white hole horizon and right boundary as the black hole. This symmetry is also reflected in the correlation plots in Fig. 6, where we can observe both horizons radiating away identically. Also, the only instability in this circuit is due to the horizons, leading to a much slower collapse.
VI Perspective on wormhole evaporation over long times

### v4

@@ 3 Setup
Crucially, the nearest-neighbour model provides an intriguing conceptual loop hole: here, the overtilt at $k\approx 0$ has no partner point with inverted tilt, very simply because the dispersion relation satisfies the periodicity constraint in $k$ by taking, in a sense, a “detour” in the complex plane – by courtesy of the EP. Consequently, the spectrum crosses the zero energy line only once. We are thus able to conclude that there is after all a case, where black and white hole horizons can be created independently in a lattice. This comes at the cost of a spontaneous instability within the overtilted region, which, importantly, is present even in the absence of an event horizon. This finding will be of great importance when interpreting the origin of radiation in the presence of event horizons in Sec. 5 below.
4 Realizing negative inductance through quench
@@
As noted above, an overtilt in the dispersion relation is realized by means of either negative capacitances or negative inductances. We here present ideas for the engineering of both features, and then argue, why we expect the latter to be more feasible, thus demonstrating claim (ii) in Sec. 2.2,
@@
Negative (and other nonlinear) capacitances are an interesting topic within quantum circuits, which date back to various pioneering ideas by Little [128] and Landauer [129], and have recently been a focal point of research, either in the form of ferroelectric materials [130, 131, 132, 133], or capacitively-coupled polarizers inducing pairing in quantum dots [134, 135]. Specifically within the circuit QED context, it has recently been proposed that negative capacitances exist in the sense of an emf-induced renormalization [60], or that the charge in the quantum phase slip energy contribution can be fractionalized and thus rendered incommensurate [136, 137]. There is, however, the problem that negative capacitances correspond to an electrostatic instability, and are thus only meaningful either as partial capacitances (where another capacitance in parallel guarantees total positivity on a given charge island) or in a transient regime. Since we are interested in the latter (after all, Hawking radiation is manifestly a transient feature), this means that the capacitance would have to be tunable on very fast time scales, which seems challenging, especially on the small scales of the here proposed circuit QED architecture.
@@
This is why we move on to ideas on how to bring about a transient behaviour within the inductive (Josephson junction) element. Remember that we focus on a regime of $E_{J}>E_{C}$ where the cosine of the Josephson energy can be approximated as a parabola with the inductive energy $E_{L}\sim E_{J}$ of an effectively regular linear inductor (rendering the resulting plasmon field effectively non-interacting, as shown above). Nonetheless, we can exploit the nonlinear inductive behaviour. Suppose there is a mechanism, allowing for a fast, transient switch from $\cos(\phi_{j}-\phi_{j-1})$ (e.g., for the nearest neighbor model) to $\cos(\phi_{j}-\phi_{j-1})$. If that switch occurs sufficiently fast, the circuit degrees of freedom have no time to adjust, such that the many-body wave function immediately after the switch remains localized on what is now the maximum (and not the minimum) of the cosine. Expanding for small phase differences localized around said maximum, we can realize a sign switch of the inductive energy, $E_{L}\rightarrow-E_{L}$, leading to the sought-after overtilt in the dispersion relation (see the orange lines in Fig. 3). Specifically, in the next section, we discuss the possibility of applying an inverted inductance only in part of the system to create event horizons separating regions with $u<|v|$ from unquenched regions with $u>|v|$.
@@
Figure 5: Illustration of non-local effects of the current loop flux drive. In order to create the target phase shift profile of exactly $\pi$ within a connected region in the junction array (from $j=j_{0}$ to $j=j_{1}$), the corresponding applied current profile as a function of site index $j$ needs to be slightly nonlocal. We have plotted three current profiles, that were obtained for a system with 50 sites with $j_{0}=15$ and $j_{1}=35$, with different ratios of $\Delta l$ to $\Delta x$ as shown in the graph. The parameter $R$ does not have a significant effect on the nature of the current profile, only on the magnitude of the ration $I_{j}/I_{0}$, hence we set $R/\Delta x=1$.
@@
$\phi^{\text{ext}}_{j}=\frac{2\pi}{\Phi_{0}}\int_{\mathcal{L}_{j}}d\mathbf{l}\cdot\mathbf{A}\ .$ (22)
@@
For the matrix given in Eq. (24), we can explicitly compute a required current profile that needs to be applied in order to swap the phases by $\pm\pi$ within a connected part of the 1D chain. Blue dots in Fig. 5 show that if $\Delta l\ll\Delta x$, $\alpha$ is almost perfectly diagonal, such that there is a one-to-one correspondence between the desired target phase profile and the required current profile that has to be applied to the individual loops. Note however that the onset of non-local effects may play a role even for (moderately) small $\Delta l/\Delta x$ (orange and brown dots in Fig. 5), where we need to apply a slightly nonlocal current profile to have a sharp, local switch from positive to negative inductances within the 1D chain.
Note that the flux drive may in general not only couple to the Josephson junctions but also to the gyrators shunted in parallel. In accordance with Ref. [60], such a stray coupling can be neglected if the surface charges induced by the time-dependent flux drive (these are the charges screening the induced electric field) are mostly allocated near the junction. We expect this scenario to be plausible if junction and gyrator are sufficiently spatially separated with respect to the length scale on which the magnetic field varies. If such a spatial separation cannot be guaranteed, then one would have to take into account the microscopic and geometric details of the gyrator which is beyond the scope of this work.
@@
To conclude this section, we point out similarities and differences to previous proposals creating analog event horizons and Hawking radiation in solid state systems. Specifically, the analog Hawking radiation previously predicted in Josephson junction arrays [33] relies on the creation of solitons, where within the finite extension of the soliton (over many lattice sites), the inductance can likewise be regarded as effectively negative within the soliton profile. The procedure we produce on the other hand involves creating $\pm\pi$ shifts on a connected series of junctions, which in essence corresponds to a tightly packed generation of “half”-solitons with the width of a single lattice site. Our system thus seems arguably much less stable. But first, as indicated already, the nonreciprocity provided by the gyrators allows for a stabilization of the system even with inverted inductors, due to an extra (positive) inductive contribution. Second, the extra tilting mechanism provided by the gyrators allows for the explicit realization of arbitrary (on-demand) spacetime geometries, and in addition, the explicit creation of wormholes with (in principle) distinguishable black and white hole horizons. As for the evolution of the unstable system beyond transient times, we provide some thoughts in Sec. 6. Another important difference is that we here have in principle perfect control on the position of the horizon, contrary to the solitons studied in Ref. [32, 33], which are autonomously moving with respect to the lab frame. Moreover, we can create an event horizon with the precision of a single lattice site. This is in contrast to all other analog Hawking radiation proposals (cold atoms, circuits, tilted Weyl semimetals), where the analog metric always changes over a finite width (a healing length). This is an interesting caveat specifically regarding the existence of a well-defined Hawking temperature, or more generally, how to make sense of horizons with diverging surface gravity, as already indicated in Sec. 2.
5 Analog horizons via flux quench
@@
Once the optimization problem of the current profile has been solved, the sign of the Josephson energy can be flipped for a region of the circuits described by the Hamiltonians (15) and (16), resulting in two horizons in the circuits (Fig. 2c).
@@
The existence of tilted dispersion relations and horizons can be probed by means of injecting wave packets within the chains, and probe their time of flight. For example, within an overtilted region, wave packets can only move in one direction. At a white hole horizon, an incoming wave packet comes to a stand still. Such features represent a first experimentally available signature of the nontrivial spacetime geometry. Note that for the time evolution of the wave packets, the distinction between nearest and next-to-nearest neighbour coupling is not that important, under the condition that the experimenter is able to prepare wave packets close to $k\approx 0$ or $k\approx\pi$. These two types of excitations clearly distinguish between black and white hole horizon as they probe only a local part of $k$ space. In addition, aspects of the wormhole stability (or instability) are not immediately visible in the wave packet amplitude (until the moment when the evaporation starts changing the spacetime geometry itself, see also Sec. 6 below).
@@
We therefore present in addition a more sophisticated measurement which is able to probe radiation due to instabilities in a more direct fashion. Namely, this section examines how to characterize the system via two point charge and phase correlation functions. As foreshadowed above, while for a stable system, quantum fluctuations remain bounded over time, they start to diverge in the presence of complex eigenvalues. In the quadratic approximation, see Eq. (3), it is convenient to express the charge and phase operators for the circuits (15) and (16), in terms of bosonic operators
@@
that obey the following commutation relations $a_{j},a^{\dagger}_{j^{\prime}}]=\delta_{jj^{\prime}}$, $a_{j},a_{j^{\prime}}]=0$. Now, calculating the two point correlation functions for charge and phase operators turns into calculating the correlation functions of the form $\langle a_{j}(t)a_{j^{\prime}}^{\dagger}(t^{\prime})\rangle$, $\langle a_{j}^{\dagger}(t)a_{j^{\prime}}^{\dagger}(t^{\prime})\rangle$, $\langle a_{j}(t)a_{j^{\prime}}(t^{\prime})\rangle$, and $\langle a_{j}^{\dagger}(t)a_{j^{\prime}}(t^{\prime})\rangle$.
We are going to calculate these correlations using two methods: 1) direct diagonalization and 2) an extension of Klich’s determinant formula [141]. The reason for the multi-pronged approach will be explained in detail below. To summarize, 1) is in principle very straightforward to implement and program, and is more versatile, as it would allow computing all types of observables, not only correlations. But diagonalization is not applicable in the presence of EPs, which is where method 2) comes into play, since it only requires exponentiation of the matrix with EPs.
As previously stated, once the system has been quenched, its state is in a highly excited state, were the new ground state is removed from the theory (a well-defined procedure for times sufficiently close to the quench). In order to anchor the theory to a well-defined ground state (necessary to compute observables in a well-defined way), we choose to perform computations with respect to the ground state of the unquenched system. Since the diagonalization of the quenched Hamiltonian is being performed with respect to a state that is in general not its eigenstate, we will need a generalised version of bosonic Bogoliubov transformations. These transformations are discussed in Appendix B. We find that in presence of horizons, both kind of circuits: one with nearest neighbor coupling via JJs and the other one with only next nearest neighbor coupling, give complex eigenvalues for the Bogoliubov transformations.
Implementing the Bogoliubov transformations on the quenched Hamiltonian, mentioned in the previous paragraph, requires diagonalizing a non-Hermitian matrix. This can pose a problem if the matrix has defective eigenvalues (EPs). We therefore suggest another method of calculating the correlations after the quench. Consider the following way to write a two point correlation
$\left\langle a^{\dagger}_{i}(t)a_{j}(t)\right\rangle=\partial_{\chi}\left\langle e^{\chi a^{\dagger}_{i}(t)a_{j}(t)}\right\rangle_{\chi=0}=\lim_{\beta\to\infty}\frac{\text{Tr}\left(e^{i\mathcal{H}_{>}t}e^{\chi a^{\dagger}_{i}a_{j}}e^{-i\mathcal{H}_{>}t}e^{-\beta\mathcal{H}_{<}}\right)}{\text{Tr}\left(e^{-\beta\mathcal{H}_{<}}\right)},$ (27)
@@
The system we will study here has hundred sites ($J=100$) with the wormhole located between $j_{0}=30$ and $j_{1}=70$. For convenience, we impose periodic boundaries in position space (i.e., the 1D chain is closed into a loop). The parameters of the system are $|E^{\prime}_{L}|=|E_{L}|=3.42E_{C}$, $G=2.4$ and $M=3*10^{-4}E_{C}$. As already pointed out, in each of the following figures of correlation functions the result could be obtained by both diagonalization and the generalized Klich determinant. The time steps in the following plots are $\delta t=0.0031E_{C}^{-1}$ for the circuit with nearest neighbor coupling and $\delta t=0.031E_{C}^{-1}$ for the next-to-nearest neighbor coupling.
@@
First, let us examine the plots of correlation functions for the circuit with only nearest neighbor coupling, Fig. 6. We see that both phase and charge quantum fluctuations diverge after the quench. The system has two boundaries (red dotted lines) between the wormhole and the normal regions, that act as the horizons (located at $j_{0}$ and $j_{1}$). If we initialize low energy excitations (for this circuit this means excitations near $k=0$) in the region with overtilted dispersion, they move from the left horizon to the right horizon, hence we label them as the black and the white hole horizon respectively. The black hole horizon radiates away with time, as can be inferred by the accumulation of quantum fluctuations near it with time (which indicates presence of radiation), while the white hole horizon does not radiate. This explicitly confirms the previously discussed fact that for the nearest neighbour system, black and white hole horizons are distinguishable. The wormhole region also starts radiating due to the presence of the EP (as likewise already discussed in Sec. 3). We observe that the decay in the wormhole interior is much faster than the one induced by the presence of the horizons, which leads to the correlation functions growing much more rapidly inside the wormhole, before the black hole horizon show any appreciable decay. Overall, the above thus illustrates the possibility to create lattice simulations of wormholes with distinguishable black and white hole horizons, at the cost of a strong instability of the wormhole interior.
@@
Now, let us focus on the circuit where Josephson junctions connect next nearest neighbors. Low-energy excitations in this circuit can be either near $k=0$ or $k=\pm\pi$. For excitations (inside the wormhole region) near $k=0$, similar to the previous circuit, the left boundary ($j_{0}$) acts as the black hole horizon and the right one ($j_{1}$) as the white hole. In contrast, for excitations near $k=\pm\pi$ the left boundary acts as the white hole horizon and right boundary as the black hole, hence we also refer to these as grey hole horizons. This symmetry is also reflected in the correlation plots in Fig. 7, where we can observe both horizons radiating away identically. Also, the only instability in this circuit is due to the horizons, leading to a much slower collapse.
Figure 6: System with Josephson junctions connecting the nearest neighbors, (left) time evolution of quantum fluctuations of phase difference, (right) time evolution of quantum fluctuations of conjugate charge. Red dotted lines denote the position of the apparent horizons, i.e., boundaries between the wormhole and normal regions. Here, the fluctuations visibly distinguish between a pure black (pure white) hole horizon, where quantum fluctuations accumulate (or not). More over, the interior between the two horizons is here unstable, such that quantum fluctuations diverge immediately within the entire wormhole region.
Figure 7: System with Josephson junctions connecting the next-nearest neighbors, (left) time evolution of quantum fluctuations of phase difference, (right) time evolution of quantum fluctuations of conjugate charge. Red dotted lines denote the position of the horizons. Here, (and contrary to Fig. 6) both horizons have black as well as white hole character, such that the two horizons are not easily distinguised when considering the spatial dependence of the fluctuations. On the other hand, the wormhole interior is here (at least initially) stable, and quantum fluctuations grow only from the horizons outwards.
We can contrast above findings with a numerical calculation on a lattice with increasingly smooth variation of circuit parameters. Naively one would expect that there exists a continuum limit where the lattice theory maps back to the continuous Lagrangian (see Eq. (1)), where stable solutions exist (non-complex eigenvalues) and where horizons are unambiguously either black or white holes. Our calculation reveals the following (see Appendix E): while solutions with real-valued eigenfrequencies exist (reproducing thermal Hawking radiation) there remain two caveats. First of all the black/white hole ambiguity remains for the next nearest coupling model – giving rise to radiation close to both $k=0$ and $k=\pi$ sectors. Secondly, the Boltzmann factor (expressing Hawking radiation) receives a renormalisation that depends strongly on the details of the discretisation, which likewise cannot be removed by simply letting the lattice constant go to zero. Overall we can distinguish three types of field theories, which are all demonstrably distinct from one another. Lattice theories can either be considered in an unstable regime where parameters change abruptly with respect to the lattice constant, or a stable regime with smooth changes. However even the latter does not simply map to the continuous model.
6 Perspective on wormhole evaporation over long times